# Counting Hallucinations in Diffusion Models

### v1

@@ 5.1 Implementation Details
This section outlines the experimental setup for both training and inference. All experiments are conducted on a cluster with 8 NVIDIA A100 GPUs for parallel training and inference.
Training of Diffusion Models. We train a DDPM (Ho et al., 2020) from scratch for the ToyShape dataset and fine-tune latent diffusion models (LDMs) (Rombach et al., 2022), pretrained on the CelebA-HQ dataset (Karras et al., 2018), for the SimObject and RealHand datasets. The time step $T$ is set to $1,000$ with a standard linear noise schedule. Images are resized to 128×128 for DDPMs and 256×256 and LDMs. We train the models using the Adam optimizer (Kingma, 2015) with a learning rate of 0.0001, running for 150k, 300k, and 80k steps on the ToyShape, SimObject, and RealHand datasets, respectively, with an effective batch size of 256.
Inference of Diffusion Models. For the ancestral sampling (i.e., DDPM), the sampling steps are set to 1,000, matching the diffusion steps during the training process. For the ODE-based solvers, we consider both first-order and second-order methods: DPM-Solver-1 (equivalent to DDIM (Song et al., 2021)) as a first-order solver, and DPM-Solver-2 (Lu et al., 2022) as a second-order solver. Each solver is evaluated under three widely-used denoising step configurations: 25, 50, and 100 steps. For the RealHand dataset, the counting model employs a YOLO detector with a confidence threshold of 0.3 and an IoU threshold of 0.1. Unless noted otherwise, the number of generated samples matches the number of training samples in each evaluation, and all experimental results are averaged over three different random seeds.
@@ 5.2 Quantifying Counting Hallucinations under Different Denoising Conditions
Table 1 presents quantitative results on counting hallucinations under varying denoising conditions, including solver types, ODE solver orders, and initial noise settings across three datasets: ToyShape, SimObject, and RealHand. Key observations are as follows.
(1) Increasing sampling steps in ODE solvers reduces counting hallucinations in synthetic datasets but exacerbates them in RealHand. As shown in Table 1, under commonly used ODE solver settings (25, 50, and 100 steps), increasing the number of sampling steps generally decreases the counting hallucination rate (CHR) in ToyShape and SimObject, but tends to increase it in RealHand, except at 100 steps with DPM-Solver-2. This contrasting behavior suggests that synthetic datasets benefit from finer solver granularity due to their simpler and more structured distributions, whereas real-world datasets such as RealHand exhibit more complex distributions, where additional steps may overfit local inconsistencies and thereby amplify hallucinations.
(2) Employing higher-order ODE solvers effectively mitigates non-counting failures but exacerbates counting hallucinations in RealHand. As shown in Table 1, comparing DPM-Solver-1 and DPM-Solver-2 on the RealHand dataset reveals that the latter reduces the non-counting failure rate but increases the counting hallucination rate. Nevertheless, this strategy consistently lowers the overall failure rates across datasets. For example, reducing CHR in synthetic datasets and TFR in RealHand, despite the increase in CHR observed in RealHand. These results underscore the need to develop ODE solvers that can jointly minimize non-counting failures while maintaining control over counting hallucinations in practical scenarios.
(3) DDPM substantially outperforms ODE solvers in mitigating both counting hallucinations and non-counting failures. Across all comparisons, ancestral sampling (i.e., DDPM) consistently achieves the lowest CHR, NCFR, and TFR, indicating that DDPM provides a practical lower bound for quantifying the failure rates of generative models. These results suggest that ancestral sampling serves as an effective strategy for minimizing hallucinations, particularly when computational efficiency is not the primary concern.
Table 1: Counting hallucination rate (CHR; the number of counting hallucinations divided by total generated samples) are reported under different solver configurations and initial noise settings across three datasets: ToyShape, SimObject, RealHand. For RealHand, we also report non-counting failure rate (NCFR) capturing non-counting failure samples. The total failure rate (TFR) is defined as the sum of counting hallucination rate (CHR) and non-counting failure rate (NCFR). Diffused and Normal refer to the ground-truth initial noise and standard Gaussian noise, respectively.
@@
(5) Greater morphological complexity in countable objects increases the likelihood of counting hallucinations. Comparing the counting hallucination rates (CHR) across ToyShape, SimObject, and RealHand reveals that diffusion models are increasingly prone to counting hallucinations as object morphology becomes more complex, from simple geometric forms (e.g., triangle), to moderately complex natural objects (e.g., apple), to articulated biological structures (e.g., human finger). This trend arises because higher morphological complexity introduces intricate spatial relationships and fine-grained details that challenge the model’s ability to preserve accurate object counts. Consequently, the probability of erroneously adding or omitting object instances increases with structural complexity.
@@
5.3 Correlation Studies: Counting Hallucination Rate vs. FID and Non-Counting Failure Rate vs. FID
This section examines the statistical correlations between the counting-based hallucination rate and the Fréchet Inception Distance (FID), as well as between the non-counting failure rate and FID, one of the most widely used evaluation metrics in image generation. Since counting hallucination represents a form of factualness hallucination, it is natural to ask: Does a better perception visual quality, indicated by a lower FID, necessarily imply fewer counting hallucinations by reflecting a smaller distribution gap between generated and training data? The answer is no necessarily. The following evidence supports this conclusion, with each paragraph presenting one piece of evidence.
The correlations between CHR and FID are dataset-dependent, rather than intrinsic. From Table 2, the Pearson correlation between counting hallucination rate (CHR) and FID is strongly positive in SimObject ($r$ > 0.87), indicating that lower FID corresponds to less counting hallucinations. In contrast, RealHand exhibits a strong negative correlation ($r$ = –0.9134), suggesting the opposite trend. Figure 3 directly illustrates this negative correlation.
The correlations between CHR and FID are solver-dependent, rather than intrinsic. As shown in Table 2, including DDPM (ancestral sampler) results reduces the consistency of the observed correlations between CHR and FID. For example, the Spearman coefficient $\rho$ on SimObject decreases from 0.8452 to 0.7538, whereas on RealHand it drops sharply from -0.8286 to -0.1429.
@@
The correlations between NCFR and FID are consistently strong and positive across different solver types and sampling conditions. In contrast to the variable relationship observed between counting hallucination rate (CHR) and FID, the association between NCFR and FID is direct and robust. Specifically, whether or not DDPM results are included, both the Pearson ($r$) and Spearman ($\rho$) coefficients exceed 0.94, with very small p-values, indicating high statistical significance.
(a) Correlation trend on DPM-Solver-1.
(b) Correlation trend on DPM-Solver-2.
Figure 3: Correlation trends between FID and three failure rate metrics on RealHand, including counting hallucination rate (CHR), non-counting failure rate (NCFR), and totle failure rate (TFR), across different ODE solvers: (a) DPM-Solver-1 and (b) DPM-Solver-2.
Table 2: Correlation of counting hallucination rate (CHR), non-counting failure rate (NCFR), and total failure rate (TFR) with FID. ‘’incl.’ denotes DDPM results integrated with those from various ODE solvers across sampling conditions (e.g., initial noise, sampling steps, and solver orders).
CHR vs. FID NCFR vs. FID TFR vs. FID
Correlation measure SimObject RealHand RealHand RealHand
$r$ or $\rho$ p-value $r$ or $\rho$ p-value $r$ or $\rho$ p-value $r$ or $\rho$ p-value
Pearson 0.8762 0.0119 -0.9134 0.0109 0.9977 <<0.0001 0.9963 <0.0001
Pearson (incl. DDPM) 0.4925 0.1482 0.0602 0.8980 0.9902 0.0001 0.9017 0.0055
Spearman 0.8452 0.0041 -0.8286 0.0416 0.9429 0.0048 0.9999 <0.0001
Spearman (incl. DDPM) 0.7538 0.0118 -0.1429 0.7599 0.9643 0.0005 0.9999 <0.0001
@@ 5.4 How to Reduce Counting Hallucinations?
In recent years, significant improvements have been observed in the perceptual quality of images generated by diffusion models, as reflected in metrics such as FID. However, counting hallucinations remain prevalent. This may be because conventional metrics, including FID, primarily capture global perceptual realism and do not reflect semantic inconsistencies, such as incorrect object counts. As shown in Table 1 and the correlation analyses in Sec. 5.3, commonly applied numerical strategies, such as employing solvers with higher orders or increasing sampling steps, effectively reduce total failure rates but fail to address counting hallucinations in generating real-world images.
@@
Interestingly, the lower counting hallucination rates observed in simpler countable objects (as shown in Observation 5 in 5.2) suggest that diffusion models can maintain accurate object counts when the underlying visual structure is sufficiently simple, such as in masked objects. This finding raises an important question: if diffusion models jointly denoise global visual features and factual constraints (e.g., object counts) within a shared latent space (which we refer to as joint-diffusion models (JDM)), could counting hallucinations be effectively reduced?
Specifically, we perform a channel-wise concatenation of the hand masks, which are obtained using SAM-2 (Ravi et al., 2024), to provide explicit structural constraints to the diffusion model. This mechanism directly steers the generative process at the pixel level, enforcing adherence to a predefined spatial layout. While the structural constraint is explicit, higher-level semantics, such as correct finger counts and plausible hand poses, emerge implicitly. By jointly learning visual representations and structural constraints within a shared latent space, the proposed JDM enforces semantic consistency and visual plausibility. As shown in Table 3, JDM consistently achieves lower failure rates compared to latent diffusion models (LDM), including CHR and NCFR, across diverse sampling conditions, highlighting its effectiveness in accurately generating structurally intricate objects that are susceptible to counting ambiguities.
Figure 4: Diagram of the proposed joint-diffusion model (JDM), which integrates hand segmentation masks as pixel-level structural constraints. By jointly learning visual representations and structural constraints within a shared latent space, JDM enforces high-level semantic consistency, leading to anatomically plausible hand poses and correct finger counts.
@@
Table 3: Comparison of failure rates between the LDM and our proposed JDM on the RealHand dataset across different ODE solvers and sampling steps. Lower values indicate better performance.
@@
Counting Hallucination Rate (CHR) $\downarrow$ Non-counting Failure Rate (NCFR) $\downarrow$
Solver Name DPM-Solver-1 DPM-Solver-2 DPM-Solver-1 DPM-Solver-2
Sampling Steps 25 50 100 25 50 100 25 50 100 25 50 100
LDM (Rombach et al., 2022) 12.95 13.85 14.55 14.48 15.99 15.43 18.06 12.51 10.63 9.33 7.22 8.94
JDM (ours) 11.41 10.67 10.51 10.12 9.66 10.07 2.94 2.42 2.29 3.51 4.63 2.26
@@ 6 Conclusion
In this work, we focus on a specific type of hallucination in diffusion models: counting hallucination. To systematically quantify this phenomenon, we introduce CountHalluSet, a dataset suite consisting of three datasets and a practical methodology for measuring counting hallucinations in image generation. Our quantification experiments yield several key findings. For example, commonly adopted numerical techniques in diffusion models help mitigate counting hallucinations in synthetic datasets but aggravate them in the RealHand dataset. Furthermore, our correlation analyses reveal that widely used perceptual quality metrics such as FID fail to capture the severity of counting hallucinations consistently.

### v2

@@ 5.1 Implementation Details
Training of Diffusion Models. We train a DDPM (ddpm) from scratch for the ToyShape dataset and fine-tune latent diffusion models (LDMs) (ldm), pretrained on the CelebA-HQ dataset (karras2018celebahq), for the SimObject and RealHand datasets. The time step $T$ is set to $1,000$ with a standard linear noise schedule. Images are resized to 128×128 for DDPMs and 256×256 and LDMs. We train the models using the Adam optimizer (adam) with a learning rate of 0.0001, running for 150k, 300k, and 80k steps on the ToyShape, SimObject, and RealHand datasets, respectively, with an effective batch size of 256.
Inference of Diffusion Models. For the ancestral sampling (i.e., DDPM), the sampling steps are set to 1,000. For the ODE-based solvers, we consider both first-order and second-order methods: DPM-Solver-1 (equivalent to DDIM (ddim)) as a first-order solver, and DPM-Solver-2 (dpm-solver) as a second-order solver. Each solver is evaluated under three widely-used denoising step configurations: 25, 50, and 100 steps. For the RealHand dataset, the counting model employs a YOLO detector with a confidence threshold of 0.3 and an IoU threshold of 0.1. Unless noted otherwise, the number of generated samples matches the number of training samples in each evaluation, and all experimental results are averaged over three different random seeds.
@@ 5.2 Quantifying Counting Hallucinations under Different Denoising Conditions
Table 1 presents quantitative results on counting hallucination rate (CHR), non-counting failure rate (NCFR), and total failure rate (TFR) under varying denoising conditions, including solver types, ODE solver orders, and initial noise settings across three datasets: ToyShape, SimObject, and RealHand. Several key observations are as follows.
(1) Increasing sampling steps reduces counting hallucinations in synthetic datasets but exacerbates them in RealHand. As shown in Table 1, under commonly used ODE solver settings (25, 50, and 100 steps), increasing the number of sampling steps generally decreases counting hallucination rates in ToyShape and SimObject, but tends to increase it in RealHand, except at 100 steps with DPM-Solver-2. This contrasting behavior suggests that synthetic datasets benefit from finer solver granularity due to their simpler and more structured distributions, whereas real-world datasets such as RealHand exhibit more complex distributions, where additional steps may overfit local inconsistencies and thereby amplify hallucinations.
(2) Employing higher-order ODE solvers effectively mitigates non-counting failures but exacerbates counting hallucinations in RealHand under the standard inference steps. As shown in Table 1, comparing DPM-Solver-1 and DPM-Solver-2 on the RealHand dataset at the standard sampling steps (e.g., 25 and 50 steps) reveals a consistent trade-off: while the higher-order solver substantially reduces non-counting failures, it simultaneously leads to a higher counting hallucination rate. A similar trend emerges when increasing the number of sampling steps for both DPM-Solver-1 and DPM-Solver-2. Overall, although using more computation (i.e., “better” sampling conditions) consistently improves overall perceptual quality (i.e., lower non-counting failure rates), it does not enforce factual correctness of finger counts.
(3) DDPM substantially outperforms ODE solvers in mitigating both counting hallucinations and non-counting failures. Across all comparisons, ancestral sampling (i.e., DDPM) consistently achieves the lowest CHR, NCFR, and TFR, indicating that DDPM provides a practical lower bound for quantifying failure rates of generative models. These results suggest that ancestral sampling serves as an effective strategy for minimizing hallucinations.
Table 1: Counting hallucination rates (CHR; the number of counting hallucinations divided by total generated samples) are reported under different solver configurations and initial noise settings across three datasets: ToyShape, SimObject, RealHand. For RealHand, we also report non-counting failure rate (NCFR), which measures failure cases captured by the counting-ready indicator. The total failure rate (TFR) is defined as the sum of CHR and NCFR. Diffused and Normal refer to the ground-truth initial noise and standard Gaussian noise, respectively.
@@
(5) Greater morphological complexity in countable objects increases the likelihood of counting hallucinations. Comparing CHR across ToyShape, SimObject, and RealHand indicates that diffusion models tend to exhibit higher CHR as the morphological complexity of the target objects increases. This suggests that fine-grained details and complex spatial relationships inherent to morphologically complex objects pose additional challenges for diffusion models to maintain accurate object counts, leading to more frequent additions or omissions.
@@
5.3 Correlation Analysis Between Counting Hallucination Rate, Non-Counting Failure Rate, and Perceptual Metrics
Counting hallucination is a form of factualness hallucination, so it is natural to ask: Does a better perceptual metric, such as a lower FID, necessarily imply fewer counting hallucinations by reflecting a smaller distribution gap between generated and training data? The answer is not necessarily.
To examine this relationship, we perform correlation analysis between CHR and perceptual metrics, as well as between the NCFR and the same metrics. We focus on two widely used perceptual measures: a distribution-based metric, FID (fid), and an instance-based metric, MUSIQ (ke2021musiq). For correlation estimation, we adopt three complementary measures: Pearson correlation (lee1988pearson) (linear dependence), Spearman rank correlation (spearman1904proof) (monotonic dependence), and Distance Correlation (szekely2007distcorre) (general nonlinear dependence). The following evidence supports our conclusion.
The correlations between CHR and perceptual metrics are dataset-dependent and solver-dependent rather than intrinsic. As shown in the first and third rows of Table 2, the Pearson correlations between CHR and FID vary substantially across datasets (i.e., positive for SimObject, negative for RealHand, and positive again for PTI-Hand), while the corresponding Spearman correlations shift from positive to negative to negative. Distance correlations also decrease from 0.82 (SimObject) to 0.74 (RealHand) to 0.47 (PTI-Hand), suggesting that more complex generative content weakens the association between CHR and FID. Moreover, integrating DDPM results with those from ODE solvers further attenuates these correlations.
@@
The correlations between NCFR and perceptual metrics are consistently strong and positive. In contrast to the weak correlations observed between CHR and perceptual metrics, the correlations between NCFR and perceptual metrics is strong and robust across all settings. For example, all correlation coefficients are over 0.88 in NCFR vs. FID, with very small p-values, indicating high statistical significance.
Table 2: Correlation of counting hallucination rate (CHR), non-counting failure rate (NCFR), and total failure rate (TFR) with perceptual metrics, including FID and MUSIQ. Results for PTI-Hand are obtained by prompting a pretrained Stable-Diffusion-3.5-Medium model with ten hand-generation prompts to produce 5050 samples, matching the size of the RealHand dataset. FID for PTI-Hand is computed between RealHand images and the model-generated samples. Here, $p$ denotes the p-value assessing the statistical significance of the correlation $r$ or $\rho$. $\dagger$} means DDPM results integrated with those from various ODE solvers across sampling conditions.
CHR vs. FID NCFR vs. FID CHR vs. MUSIQ NCFR vs. MUSIQ
Metric SimObject RealHand PTI-Hand RealHand PTI-Hand RealHand PTI-Hand RealHand PTI-Hand
$r$ / $\rho$ / $r_{d}$ $p$ $r$ / $\rho$ / $r_{d}$ $p$ $r$ / $\rho$ / $r_{d}$ $p$ $r$ / $\rho$ / $r_{d}$ $p$ $r$ / $\rho$ / $r_{d}$ $p$ $r$ / $\rho$ / $r_{d}$ $p$ $r$ / $\rho$ / $r_{d}$ $p$ $r$ / $\rho$ / $r_{d}$ $p$ $r$ / $\rho$ / $r_{d}$ $p$
Pearson ($r$) 0.87 0.01 -0.97 <<0.01 0.20 0.69 0.99 <<0.01 0.88 <<0.01 0.93 <<0.01 -0.02 0.95 -0.97 <<0.01 -0.99 <<0.01
$\text{Pearson}^{{\dagger}}$ ($r$) 0.49 0.14 -0.53 0.13 0.96 <<0.01 0.46 0.20 -0.98 <<0.01
Spearman ($\rho$) 0.84 <<0.01 -0.97 <<0.01 -0.14 0.78 0.99 <<0.01 0.92 <<0.01 0.80 0.01 0.42 0.33 -0.88 <<0.01 -0.46 0.29
$\text{Spearman}^{{\dagger}}$ ($\rho$) 0.75 0.01 -0.51 0.15 0.95 <<0.01 0.40 0.28 -0.91 <<0.01
Distance Correlation ($r_{d}$) 0.89 0.97 0.47 0.99 0.91 0.93 0.42 0.97 0.99
$\text{Distance Correlation}^{{\dagger}}$ ($r_{d}$) 0.82 0.74 0.95 0.71 0.97
(a) O1: FR vs. Steps
(b) O1: Quality vs. Steps
(c) O2: FR vs. Steps
(d) O2: Quality vs. Steps
(e) 25 steps: FR vs. Order
(f) 25 steps: Quality vs. Order
Figure 3: Comparison of trend consistency between DM-RealHand (ours) and PTI-Hand (pretrained Stable-Diffusion-3.5-Medium). O1 and O2 denote first- and second-order ODE solvers (DPM-solver-1 / DPM-solver-2 for DM-RealHand; Euler / Heun for PTI-Hand). FR metrics include CHR and NCFR; quality metrics include FID and MUSIQ. Across variations in sampling steps and solver order, both models exhibit highly consistent trends.
5.4 Generalization of RealHand Findings to Pretrained Text-to-Image Models
In this section, we evaluate counting hallucinations in a large-scale pretrained text-to-image model. Specifically, we assess Stable-Diffusion-3.5-Medium (sd_v3) using the counting-ready indicator and counting model introduced in Sec. 4.2.3. More details are provided in Appendix A.8. As shown in Figure 3, the way solver type and sampling steps affect counting hallucinations in large pretrained diffusion models aligns closely with the trends revealed by our controlled RealHand experiments. This consistency highlights the validity of the findings presented in Sec. 5.2 and underscores the broader applicability of our evaluation protocol. Correlation results in Table 2 further support this conclusion. In specific, weak correlations between CHR and perceptual metrics and strong correlations between NCFR and perceptual metrics are also observed. Together, these results demonstrate the robustness and generality of our findings and show that the proposed evaluation protocol transfers reliably to pretrained text-to-image models.
@@ 5.5 How to Reduce Counting Hallucinations?
The lower counting hallucination rates observed for simpler countable objects (see Observation 5 in Sec. 5.2) suggest that diffusion models more reliably capture counting-based factuality when the underlying visual structure exhibits low morphological complexity (e.g., masked objects). This finding raises an important question: if diffusion models were to denoise general visual representations together with explicit, simple structural constraints that are easier to model, such as object masks that encode basic topology and counts, within a shared latent space, could such a joint learning paradigm mitigate counting hallucinations? We refer to such models that jointly learn both general visual representations and structural constraints during denoising as joint-diffusion models (JDM).
Specifically, we perform a channel-wise concatenation of the latent representation of hand mask with that of the corresponding RGB image, as shown in Fig. 4. Then we finetune the UNet pretrained on CelebA-HQ (karras2018celebahq) (following the RealHand configuration described in Sec. 5.1) using the RealHand images and their associated masks, extracted using SAM-2 (ravi2024sam2). To accommodate both modalities, the input and output channels of the denoising UNet are expanded to match the combined latent dimensionality of the image and mask latents (i.e., 3+3=6 channels).
As shown in Table 3, JDM consistently achieves lower failure rates than the vanilla LDM, including CHR and NCFR, across diverse sampling conditions, highlighting its effectiveness in accurately generating structurally intricate objects that are prone to counting ambiguities. On the other hand, JDM yields higher FID. This degradation likely stems from the increased denoising burden introduced by jointly modeling two modalities under a small-scale training regime, which forces the model to trade off perceptual quality for improved structural realism. We also observe promising reductions in counting hallucinations when applying JDM to large-scale pretrained diffusion models under carefully controlled finetuning settings. More details, results and discussion can be found in Appendix A.9.
Figure 4: Diagram of the proposed joint-diffusion model (JDM). The model performs channel-wise concatenation of the latent representations of hand RGB images and their corresponding segmentation masks. This joint learning paradigm encourages anatomically plausible hand poses and promotes correct finger topology during generation, resulting in reduced CHR and NCFR.
@@
Table 3: Comparison of failure rates between LDM and our proposed JDM on the RealHand dataset across different ODE solvers and sampling steps. Lower values indicate better performance. For JDM, although the model jointly generates hand RGB images and segmentation masks, FID is computed solely on the RGB outputs to ensure comparability with baselines.
@@
Counting Hallucination Rate (CHR) $\downarrow$ Non-counting Failure Rate (NCFR) $\downarrow$ Fréchet Inception Distance (FID) $\downarrow$
Solver Name DPM-Solver-1 DPM-Solver-2 DPM-Solver-1 DPM-Solver-2 DPM-Solver-1 DPM-Solver-2
Sampling Steps 25 50 100 25 50 100 25 50 100 25 50 100 25 50 100 25 50 100
LDM 12.95 13.85 14.55 14.48 15.99 15.43 18.06 12.51 10.63 9.33 7.22 8.94 29.86 26.43 25.21 24.48 23.42 24.54
JDM (ours) 11.41 10.67 10.51 10.12 9.66 10.07 2.94 2.42 2.29 3.51 4.63 2.26 37.69 35.51 34.68 34.63 32.07 33.89
@@ 6 Conclusion
In this work, we take the first step toward systematically quantifying hallucinations in diffusion models and offer new insights into improving factual generation in this field. Focusing on counting hallucinations, we introduce CountHalluSet and a practical and reproducible methodology for measuring counting hallucinations in practice. Through comprehensive and controlled quantification experiments, we show that commonly adopted numerical strategies often fail to mitigate counting hallucinations in generating human hands. Moreover, our correlation analyses demonstrate that widely used perceptual metrics (e.g., FID and MUSIQ) are unable to consistently reflect the severity of counting hallucinations. Experiments on large-scale pretrained text-to-image models further corroborate both the generality and the validity of our findings.